OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -1 2x - y = 4

OpenStudy (anonymous):

x - 2y = -1 2x - y = 4

OpenStudy (anonymous):

Plug in a made up value for one of the variables to figure out what the other would be to make the equation valid. For example, if you set y to equal 1, then x has to equal 1, because 1 minus 2 equals -1.

OpenStudy (anonymous):

need help fast

OpenStudy (anonymous):

if y equaled 2, x would have to equal 5 to satisfy the equations parameters, and those x and y values you achieve are your coordinates.

OpenStudy (anonymous):

x=11 , y=6.

OpenStudy (anonymous):

11,6 is not the answer

OpenStudy (anonymous):

rewrite the first to x = -1+2y and inset in the second 2(-1+2y)-y=4 and solve -2+4y-y = 4 -2+3y=4 3y = 4+2 y = 6/2 y=3 x = -1+2y x=-1+2*3 x = 5

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

x = 5, y = 3

OpenStudy (anonymous):

5,3 is not the answer

OpenStudy (anonymous):

sry 6/3 y = 2 x = 3

OpenStudy (anonymous):

rewrite the first to x = -1+2y and inset in the second 2(-1+2y)-y=4 and solve -2+4y-y = 4 -2+3y=4 3y = 4+2 y = 6/3 y=2 x = -1+2y x=-1+2*2 x = 3

jhonyy9 (jhonyy9):

x-2y=-1 /2 2x-y=4 2x-4y=-2 2x-y=4 ---------- -- -3y=-6 y=-6/-3= 2 than y=2 --- 2x-2=4 --- 2x=6 --- x=6/2 =3 x=3 y=2